OpenStudy (anonymous):

find the solutions by completing the square? Please help?! .05x^2-x+1

OpenStudy (anonymous):

it may not be possible but i just need help?!

OpenStudy (anonymous):

@satellite73 can you help?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

if its not possible by completing the square im suppose to say why but im not sure where to even start here?

OpenStudy (anonymous):

you want \[.05x^2-x+1=0\] and you need to complete the square, so first step is to make the leading coefficient 1 since \(\frac{1}{.05}=\frac{100}{5}=20\) you need to multiply everything by 20 to accomplish this \[20(.05x^2-x+1)=x^2-20x+20\]

OpenStudy (anonymous):

it is always possible to complete the square

OpenStudy (anonymous):

now we want to solve \[x^2-20x+20=0\] first subtract 20 from both sides to get \[x^2-20x=-20\] and then we "complete the square" by saying "half of 20 is 10 and \(10^2=100\) and write \[(x-10)^2=-20+100\] i.e. \[(x-10)^2=80\]

OpenStudy (anonymous):

take the square root of both sides. don't forget the \(\pm\)

OpenStudy (anonymous):

get \[x-10=\pm\sqrt{80}\] and so \[x=10\pm\sqrt{80}\]

OpenStudy (anonymous):

since \(80=16\times 5\) you know \(\sqrt{80}=\sqrt{16}\times \sqrt{5}=6\sqrt{5}\) so you can also write the answer as \[x=10\pm6\sqrt{5}\]

OpenStudy (anonymous):

all steps are there, hope it is clear

OpenStudy (anonymous):

Thank you